# Parametrically enhancing sensor sensitivity at an exceptional point

### 12 Dec 2023 v2

@@ Abstract
We propose a scheme to enhance the sensitivity of Non-Hermitian optomechanical mass-sensors. The benchmark system consists of two coupled optomechanical systems where the mechanical resonators are mechanically coupled. The optical cavities are driven either by a blue or red detuned laser to produce gain and loss, respectively. Moreover, the mechanical resonators are parametrically driven through the modulation of their spring constant. For a specific strength of the optical driving field and without parametric driving, the system features an Exceptional Point (EP). Any perturbation to the mechanical frequency (dissipation) induces a splitting (shifting) of the EP, which scales as the square root of the perturbation strength, resulting in a sensitivity-factor enhancement compared with conventional optomechanical sensors. The sensitivity enhancement induced by the shifting scenario is weak as compared to the one based on the splitting phenomenon. By switching on parametric driving, the sensitivity of both sensing schemes is greatly improved, yielding to a better performance of the sensor. We have also confirmed these results through an analysis of the output spectra and the transmissions of the optical cavities. In addition to enhancing EP sensitivity, our scheme also reveals nonlinear effects on sensing under splitting and shifting scenarios. This work sheds light on new mechanisms of enhancing the sensitivity of Non-Hermitian mass sensors, paving a way to improve sensors performance for better nanoparticles or pollutants detection, and for water treatment.
Optomechanics, exceptional point, sensor
@@ I Introduction
Exceptional Points (EPs), Non-Hermitian degeneracy, are known in physical systems for their interesting counter-intuitive features and intriguing effects. Among the fascinating properties of EPs, we can mention nonreciprocity Thomas et al. (2016); Lau and Clerk (2018), topological transport phenomena Xu et al. (2023a, b), stopping light Goldzak et al. (2018), loss-induced suppression and revival of lasing, pump-induced lasing death and unidirectional invisibility (see Peng et al. (2014) and references therein). Recently, EPs have been used to enhance sensor sensitivity and versatile physical systems were proposed for this purpose including optical/photonic Hodaei et al. (2017); Chen et al. (2017), microwave Nada et al. (2018), plasmonics Park et al. (2020); Jiang et al. (2022), and mechanic/acoustic Rosa et al. (2021); Cai et al. (2022). Coupled optomechanical systems, which result from a coupling between an electromagnetic field and a mechanical object Djorwé et al. (2013a); Aspelmeyer et al. (2014), have also been used to engineer EPs owing to their ability to induce gain and losses depending on their sideband pumping mechanism. Exceptional points engineering in optomechanics have led to a number of interesting phenomena such as mass sensing improvement Djorwe et al. (2019a); Tchounda et al. (2023), topological energy transfer Chen et al. (2023); Xu et al. (2016), and nonlinearly induced collective phenomena Djorwe et al. (2018); Djorwé et al. (2020) among others.
@@
Exceptional point sensing is an approach which provides better sensitivity compared to conventional sensing schemes where the perturbation induces a shift in the frequency. The resulting shift/signal is generally weak and needs a pre-amplification process in order to be detected and analyzed Zhu et al. (2021). For the Non-Hermitian sensors, however, the signal is self-amplified owing to the abrupt singularity and the topological feature at the EP, leading to a giant sensitivity enhancement without a pre-amplification requirement of the detected signal. This new sensing scheme has recently attracted attention as a means to enhance the responsiveness of sensors, and has started to be practically implemented. For instance, exceptional point sensing has been used to monitor internal physiological states via wireless interrogation techniques Chen and El-Ganainy (2019), and for an efficient wireless power transfer Assawaworrarit et al. (2017); Song et al. (2021). Beside these interesting sensing applications probed at EPs, the implementation of EP-sensors is still in its infancy stage and it seems to rise up scepticism regarding noise amplification, which induces a degradation of the sensor’s performance Anderson et al. (2023). This noise issue can be of a fundamental nature owing to the eigenbasis collapse or of a technical nature related to the amplification mechanisms used for the realization of EPs. This question has been a bottleneck of EP-sensors, however, appropriate techniques have started to be used to handle and suppress this noise issue Kononchuk et al. (2022); Suntharalingam et al. (2023); Tuxbury et al. (2022). Therefore, tremendous research activities are still going on to tackle and to get rid of this limitation Chen and Zhao (2020); Liu et al. (2021); Wu et al. (2023); Khanbekyan (2023), and this includes nonlinear phenomena Li et al. (2023); Bai et al. (2022).
@@
The current work introduces a new mechanism to enhance the sensitivity of EP-sensors, which is based on a mechanical parametric driving effect. Our proposal consists of two coupled optomechanical systems, where the mechanical resonators are mechanically coupled. The parametric driving comes from a nonlinear modulation of the spring constants of the mechanical resonators Lemonde et al. (2016); Bothner et al. (2020). We investigated two sensing schemes. The first one is based on the splitting Djorwe et al. (2019a) of the EP and the second is related to the shifting Wang et al. (2019); Mao et al. (2023) of the EP. When parametric driving is switched on, we found that i) the sensitivity is greatly enhanced, ii) the EP is shifted towards low frequencies, and iii) the strength of the optical driving field corresponding to the EP increases. Moreover, it has been observed that the overall sensitivity enhancement in the shifting scenario is weak compared to the enhancement captured with the splitting scheme. Furthermore, the output spectra and transmissions of the optical cavities have been investigated under some system parameters. The effect of perturbations, the parametric driving impact and thermal noise on the EP were analyzed via the spectra. While the effect of the other parameters is confirmed as aforementioned, it has been found that thermal noise amplifies the output signal, leading to better sensitivity. This work shows how sensing based on the splitting of the EP outperforms sensors operating in a scenario where the EP is shifted. Furthermore, it also paves a way towards improvement of sensing through nonlinear effects in optomechanics Djorwé et al. (2013b); Leijssen et al. (2017); Djorwe et al. (2019b).
@@
The rest of the work is organized as follows. In Section II we describe the model and derive the related dynamical equations. The sensitivity schemes based on splitting and shifting of the EP are presented in Section III. The transmissions and output spectra of the optical cavities are analyzed in Section IV, while Section V concludes our work.
II Model and dynamical equations
@@
We propose a two-cavity system with mechanically coupled vibrating mechanical resonators, as shown in /cref[fig:Fig1]. One cavity is driven by a red (blue) driving field that creates loss (gain) in the system. In addition, the mechanical resonators are parametrically driven for sensitivity enhancement prospect as it will be pointed out later on. The Hamiltonian of this system ($\hbar=1$) is given by,
Figure 1: Sketch of our benchmark system. Cavity $1$ ($2$) is driven with red (blue) electromagnetic field to generate losses (gain). The sprint constants of the involved mechanical resonators are modulated to induce parametric driving Lemonde et al. (2016); Bothner et al. (2020).
@@
where,
$H_{\rm{0M}}$ $=$ $\sum_{j}\omega_{j}b^{\dagger}_{j}b_{j}+\omega_{c}^{j}a^{\dagger}_{j}a_{j}-g_{j}a^{\dagger}_{j}a_{j}(b^{\dagger}_{j}+b_{j}),$ (2)
$H_{\rm{PA}}$ $=$ $\sum_{j}\frac{\chi_{j}}{2}(b^{\dagger 2}_{j}e^{-i\phi_{d}}+b^{2}_{j}e^{-i\phi_{d}})$ (3)
$H_{\rm{int}}$ $=$ $-J_{m}(b_{1}b_{2}^{\dagger}+b_{1}^{\dagger}b_{2}),$ (4)
$H_{\rm{drive}}$ $=$ $i\sum_{j}(\rm{E_{j}}a^{\dagger}_{j}e^{-i\omega^{j}_{p}t}-\rm{E^{*}_{j}}a_{j}e^{i\omega^{j}_{p}t}).$ (5)

### 13 Jan 2024 v3

@@ Abstract
We propose a scheme to enhance the sensitivity of Non-Hermitian optomechanical mass-sensors. The benchmark system consists of two coupled optomechanical systems where the mechanical resonators are mechanically coupled. The optical cavities are driven either by a blue or red detuned laser to produce gain and loss, respectively. Moreover, the mechanical resonators are parametrically driven through the modulation of their spring constant. For a specific strength of the optical driving field and without parametric driving, the system features an Exceptional Point (EP). Any perturbation to the mechanical frequency (dissipation) induces a splitting (shifting) of the EP, which scales as the square root of the perturbation strength, resulting in a sensitivity-factor enhancement compared with conventional optomechanical sensors. The sensitivity enhancement induced by the shifting scenario is weak as compared to the one based on the splitting phenomenon. By switching on parametric driving, the sensitivity of both sensing schemes is greatly improved, yielding to a better performance of the sensor. We have also confirmed these results through an analysis of the output spectra and the transmissions of the optical cavities. In addition to enhancing EP sensitivity, our scheme also reveals nonlinear effects on sensing under splitting and shifting scenarii. This work sheds light on new mechanisms of enhancing the sensitivity of Non-Hermitian mass sensors, paving a way to improve sensors performance for better nanoparticles or pollutants detection, and for water treatment.
Optomechanics, exceptional point, sensor, non-Hermition, parametric drive
@@ I Introduction
Exceptional Points (EPs), non-Hermitian degeneracy, are known in physical systems for their interesting counter-intuitive features and intriguing effects. Among the fascinating properties of EPs, we can mention non-reciprocity [1, 2], topological transport phenomena [3, 4], stopping light [5], loss-induced suppression and revival of lasing, pump-induced lasing death and unidirectional invisibility (see [6] and references therein). Recently, EPs have been used to enhance sensor sensitivity and versatile physical systems were proposed for this purpose including optical/photonic [7, 8], microwave [9], plasmonics [10, 11], and mechanics/acoustics [12, 13]. Coupled optomechanical systems, which result from a coupling between an electromagnetic field and a mechanical object [14, 15], have also been used to engineer EPs owing to their ability to induce gain and losses depending on their sideband pumping mechanism. EPs engineering in optomechanics have led to a number of interesting phenomena such as mass sensing improvement [16, 17], topological energy transfer [18, 19], and nonlinearly induced collective phenomena [20, 21] among others.
@@
Exceptional point sensing is an approach which provides a better sensitivity compared to conventional sensing schemes, where the perturbation induces a shift in the frequency. The resulting shift/signal is generally weak and needs a pre-amplification process in order to be detected and analyzed [22]. For the non-Hermitian sensors, however, the signal is self-amplified owing to the abrupt singularity and the topological feature at the EP, leading to a giant sensitivity enhancement without a pre-amplification requirement of the detected signal. This new sensing scheme has recently attracted attention as a means to enhance the responsiveness of sensors, and has started to be practically implemented. For instance, exceptional point sensing has been used to monitor internal physiological states via wireless interrogation techniques [23], and for an efficient wireless power transfer [24, 25]. Besides these interesting sensing applications probed at EPs, the implementation of EP sensors is still in its infancy stage and it seems to rise up scepticism regarding noise amplification, which induces a degradation of sensor’s performance [26]. This noise issue can be of a fundamental nature owing to the eigenbasis collapse or of a technical nature related to the amplification mechanisms used for the realization of EPs. This question has been a bottleneck of EP-sensors, however, appropriate techniques have started to be used to handle and suppress this noise issue [27, 28, 29]. Therefore, tremendous research activities are still going on to tackle and to get rid of this limitation [30, 31, 32, 33], and this includes nonlinear phenomena [34, 35].
@@
The current work introduces a new mechanism to enhance the sensitivity of EP-sensors, which is based on a mechanical parametric driving effect. Our proposal consists of two coupled optomechanical systems, where the mechanical resonators are mechanically coupled. The parametric driving comes from a nonlinear modulation of the spring constants of the mechanical resonators [36, 37]. We investigated two sensing schemes. The first one is based on the splitting [16] of the EP, resulting from a perturbation of the mechanical resonance frequencies, and the second one is related to the shifting [38, 39] of the EP, due to a mismatch between the mechanical dissipations. When parametric driving is switched on, we found that i) the sensitivity is greatly enhanced, ii) the EP is shifted towards low frequencies, and iii) the strength of the optical driving field corresponding to the EP increases. Moreover, it has been observed that the overall sensitivity enhancement in the shifting scenario is weak compared to the enhancement captured with the splitting scheme. Furthermore, the output spectra and transmissions of the optical cavities have been investigated under some system parameters. The effect of perturbations, the parametric driving impact and thermal noise on the EP were analyzed via the spectra. While the effect of the other parameters is confirmed as aforementioned, it has been found that thermal noise amplifies the output signal, leading to better sensitivity. This work shows how sensing based on the splitting of the EP outperforms sensors operating in a scenario where the EP is shifted. Furthermore, it also paves a way towards improvement of sensing through nonlinear effects in optomechanics [40, 41, 42].
@@
The rest of the work is organized as follows. In Section II.1 we describe the model and derive the related dynamical equations. The sensitivity schemes based on splitting and shifting of the EP are presented in Section II.2. The transmissions and output spectra of the optical cavities are analyzed in Section II.3, while Section III concludes our work.
II Results
II.1 Model and dynamical equations
@@
We propose a two-cavity system with mechanically coupled vibrating mechanical resonators, as shown in Figure 1a. One cavity is driven by a red (blue) driving field that creates loss (gain) in the system. In addition, the mechanical resonators are parametrically driven for sensitivity enhancement prospect as it will be pointed out later on. The Hamiltonian of this system ($\hbar=1$) is given by the following equation:
@@
where
$H_{\rm{0M}}$ ::=:= $\sum_{j}\left(\omega_{j}b^{\dagger}_{j}b_{j}+\omega_{c}^{j}a^{\dagger}_{j}a_{j}-g_{j}a^{\dagger}_{j}a_{j}(b^{\dagger}_{j}+b_{j})\right),$ (2)
$H_{\rm{PA}}$ ::=:= $\sum_{j}\frac{\chi_{j}}{2}(b^{\dagger 2}_{j}e^{-{\rm i}\phi_{d}}+b^{2}_{j}e^{{\rm i}\phi_{d}}),$ (3)
$H_{\rm{int}}$ ::=:= $-J_{m}(b_{1}b_{2}^{\dagger}+b_{1}^{\dagger}b_{2}),$ (4)
$H_{\rm{drive}}$ ::=:= ${\rm i}\sum_{j}({\rm E}_{j}a^{\dagger}_{j}{\rm e}^{-{\rm i}\omega^{j}_{p}t}-\rm{E}^{*}_{j}a_{j}{\rm e}^{{\rm i}\omega^{j}_{p}t}).$ (5)